Consider a 3-year coupon bond with a face value of $1,000 and a coupon rate of 10 percent. Mr. Smith purchased this bond at par (i.e., he paid Pt= $1,000) when it was newly issued.  The market interest rate at the time the coupon bond was issued was 10 percent. One year from the time of the bond’s issue, he decides to sell the bond, i.e., the holding period is one year. At that time, the market interest rate has risen to 15 percent.

1.find the price that Mr. Smith will obtain when he sells the bond, Pt+1

2.Pt+1 will be greater than Pt. Is this statement TRUE, FALSE, or UNCERTAIN?
Answer #1

Face value = 1000, Pt =1000

coupon rate = 10% = 0.1

annual payment = coupon rate * face value = 0.1*1000 =100

total time = 3 yrs , 1 yrs passed

therefore time left = 2 yr

then present value at i =15%

= 100 * (P/A, 15%, 2) + 1000 * (P/F, 15%, 2 yrs)

= 100 * 1.62570 + 1000 * 0.75614

= 918.71

Pt+1 = 918.71

As Pt+1<Pt hence the statement is false